# BANG: Billion-Scale Approximate Nearest Neighbour Search using a Single GPU

### v2

@@ Abstract
Approximate Nearest Neighbour Search (ANNS) is a subroutine in algorithms routinely employed in information retrieval, pattern recognition, data mining, image processing, and beyond. Recent works have established that graph-based ANNS algorithms are practically more efficient than the other methods proposed in the literature. The growing volume and dimensionality of data necessitates designing scalable techniques for ANNS. To this end, the prior art has explored parallelizing graph-based ANNS on GPU leveraging its massive parallelism. The current state-of-the-art GPU-based ANNS algorithms either (i) require both the dataset and the generated graph index to reside entirely in the GPU memory, or (ii) they partition the dataset into small independent shards, each of which can fit in GPU memory, and perform the search on these shards on the GPU. While the first approach fails to handle large datasets due to the limited memory available on the GPU, the latter delivers poor performance on large datasets due to high data traffic over the low-bandwidth PCIe bus.
We introduce BANG, a first-of-its-kind technique for graph-based ANNS on GPU for billion-scale datasets, that cannot entirely fit in the GPU memory. BANG stands out by harnessing a compressed form of the dataset on a single GPU to perform distance computations while efficiently accessing the graph index kept on the host memory, enabling efficient ANNS on large graphs within the limited GPU memory. BANG incorporates highly-optimized GPU kernels and proceeds in phases that run concurrently on the GPU and CPU, taking advantage of their architectural specificities. Furthermore, it enables overlapping communication with computation that results in efficient data transfer between the CPU and GPU. We evaluate BANG using a single NVIDIA Ampere A100 GPU on three popular ANN benchmark datasets. BANG outperforms the state-of-the-art comprehensively. Notably, on the billion-size datasets, we achieve throughputs 40$\times$-200$\times$ more than the competing methods for a high recall value of 0.9. Additionally, BANG is the best in cost- and power-efficiency among the competing methods from the recent Billion-Scale Approximate Nearest Neighbor Search Challenge.
@@
Approximate Nearest Neighbour Search, Graph and tree search strategies, Information retrieval, Approximate search, vector similarity search, GPU, Big data.
@@ I Introduction
The $k$-Nearest-Neighbour-Search problem is to find the $k$ nearest data points to a given query point in a multidimensional dataset. As the dimensionality increases, exact search methods become increasingly inefficient. In order to evaluate the exact $k$-nearest-neighbours of a query point in a $d$-dimensional dataset having $n$ points, as a consequence of the curse of dimensionality, all the $n$ points must be examined; this takes $O(nd)$ time [24]. Therefore, it is common practice to use Approximate Nearest Neighbour Search (ANNS) when looking for nearest neighbours in order to mitigate the curse of dimensionality by sacrificing a small accuracy for speed. ANNS is employed as a sub-routine in various fields, such as computer vision, document retrieval, and recommendation systems. These applications require searching enormous datasets containing words, images, and documents embedded in a multidimensional space. Search queries on such massive multidimensional datasets are often processed in batches to meet the high throughput demands, as evidenced by recent efforts[49, 27, 21] leveraging the massive parallelism of GPUs for ANNS.
@@
Graph-based ANNS algorithms [31, 18, 25] have been shown to be generally more efficient in practice at handling large datasets. However, GPU implementations of these algorithms [35, 49, 21] require storing graph data structures in GPU memory, which limits their ability to handle large datasets. Even recent GPUs like the NVIDIA Ampere A100 having 80GB device memory, cannot accommodate the entire input data (graph index and data points). Prior solutions, such as sharding, effectively implemented by GGNN [21], result in high memory transfer costs. For example, even with the PCIe 4.0 interconnect operating at its peak theoretical transfer rate of 32 GB/s, transferring the DEEP1B dataset (consisting of a data size of 384GB and a graph index size of 260GB) from CPU to GPU would take 20 seconds, resulting in a low throughput of less than 500 QPS for 10,000 concurrent queries on the GPU. Alternatively, hashing and compression techniques, as showcased by SONG [49] and FAISS [27], can handle large volumes of data using a single GPU by reducing the data dimensionality or by compressing the vectors. However, these approaches have limitations when it comes to high recalls for massive datasets. It has been shown that it is feasible to attain high throughputs at high recalls on large datasets using multiple GPUs [21]. However, these approaches have a very high hardware cost which makes them undesirable. Thus, this paper explores an important question: Can we increase the throughput of ANNS queries without compromising their recall by using a single GPU?
@@
In this paper, we introduce BANG, a novel GPU-based ANNS method, which is designed to efficiently handle large datasets, with billions of points, that cannot entirely fit in GPU memory. BANG stands out by employing compressed data for accelerated distance computations on the GPU while keeping the graph index and the dataset on the CPU, thus enabling efficient ANNS on large datasets. CPU transfers neighbour information to the GPU on demand in every iteration of the search. Furthermore, in contrast to the previous approaches that treat search activity as one monolithic block, BANG divides ANNS activity into distinct phases, thus maximizing resource utilization and mitigating CPU-GPU data transfer bottlenecks. Specifically, the division of phases leads to three advantages: (1) executing different phases on CPU and GPU in a pipeline (2) optimizing each phase’s (GPU kernel’s) span characteristics separately to maximize parallelism on GPU (3) asynchronously prefetching required data from CPU to GPU for certain phases. Thus, the overall performance of BANG stems from several factors, such as CPU-GPU load balancing, highly optimised GPU kernels (for tasks such as distance calculations, sorting, and updating worklists), prefetching and pipelining.
@@
ANNS methods typically follow the index and search paradigm whereby data points are first processed to construct an efficient index, and then search queries are processed using this index. In this work, our focus is efficient Billion-scale ANN search on GPU utilizing an underlying graph index, and therefore we do not construct a graph index but instead utilize an existing one from prior work [25].
Our evaluations show that BANG significantly outperforms the state-of-the-art in all cases on three popular ANN benchmark datasets using a single NVIDIA A100 GPU. In particular, on the billion-scale datasets, we achieve throughputs 40$\times$-200$\times$ more than the competing methods for a high recall of 0.9. BANG uses only a single GPU but achieves high throughput; it demonstrates high cost efficiency and power efficiency among the competing methods from the recent Billion-Scale Approximate Nearest Neighbor Search Challenge [40].
@@
  • We propose a phased-execution strategy that enables pipelined-execution of ANNS steps on both CPU and GPU, optimizing CPU-GPU load balancing, asynchronous data transfer, and GPU kernel performance for maximum parallelism.
  • We present highly optimized GPU kernels for distance calculations and worklist updates in ANNS and implement prefetching and pipelining to effectively utilize CPU and GPU, reducing PCIe memory traffic and enhancing throughput without sacrificing recall.
@@ II-B Graph Index
Many top-performing ANNS algorithms are graph-based, achieving both high recall (true nearest neighbours) and high throughput, measured as queries per second (QPS). A graph-based ANNS algorithm runs on a proximity graph, which is pre-constructed over the dataset points by connecting each point with nearby points. Our BANG algorithm implements an efficient ANNS that can run on various proximity graphs such as kNN [21], HNSW [31], Vamana [25] etc. Manohar et al. [33] conducted a study on various CPU-based ANNS algorithms and showed that DiskANN [25] approach that runs on the Vamana graph index achieves superior throughput and recall on large datasets. Hence, we utilized the Vamana graph as the underlying search structure in the implementation of BANG. Vamana graph index is constructed by sharding the large dataset into smaller overlapping clusters. The technique iteratively processes individual clusters and uses RobustPrune and GreedySearch routines to construct directed sub-graphs on host RAM, which are then stored on SSD. The final graph is formed by uniting the edges in the sub-graphs. While prior techniques (that construct the graph index using the SNG property [6]) aim to reduce distances to the query point along the search path, Vamana further reduces hops to the query point with the RobustPrune property, resulting in a fast search.
@@ III ANN Search on GPU: Challenges in Handling Billion-scale Data
We propose BANG, a parallel implementation of ANNS on a single GPU. There are two main challenges in parallelizing the ANNS on a single GPU: (1) managing large graphs within the constraints of limited GPU memory, (2) extracting sufficient parallelism for ANNS to fully utilize the hardware resources. In the following subsections, we discuss these challenges in more detail.
@@ III-A Limited GPU Memory
A primary challenge when dealing with large datasets on GPUs is their massive memory footprint. Table I shows the sizes of various billion-scale ANNS datasets and their respective graph sizes. Notably, even with the recent A100 GPU’s maximum global memory capacity of 80GB, it is evident that the entire input data (base, graph, and query) cannot be accommodated within the GPU memory. To address this challenge, a recent work GGNN [21] effectively implemented sharding as a solution. However, this approach requires frequent swaps of both processed and unprocessed shards between GPU and host RAM, resulting in high memory transfer costs over the PCI interconnect. Further, these swaps may result in an idle time of CPU/GPU, and optimal GPU-CPU coordination is essential during iterative graph traversals of ANNS. Furthermore, sharding the entire data is not viable because even with the PCIe 4.0 interconnect operating at its peak theoretical transfer bandwidth of 32 GB/s, it will take an estimated 20 seconds to transfer the largest dataset from CPU to GPU and will result in a markedly low throughput of around 500 QPS as explained in Section I.
Although employing a multi-GPU setup can facilitate the effective distribution of shards across all GPUs to store the entire input data (as demonstrated with eight GPUs in GGNN [21]), this approach incurs significant hardware cost. As a result, our focus is on developing an efficient and cost-effective parallel implementation of ANNS on a single GPU.
@@
SIFT1B 128GB 260GB 80GB
DEEP1B 384GB 260GB 95GB
SPACEV1B 140GB 260GB 112GB
Alternatively, hashing and quantization techniques can handle large data on a single GPU. Hashing effectively reduces data dimensionality to make it fit within GPU memory, as demonstrated by SONG [49], which compressed a 784-byte vector to 64 bytes in the MNIST8M dataset. However, hashing is better suited for smaller datasets and may not achieve high recall on datasets with billions of points. Data compression, as demonstrated by FAISS [27], can accelerate query processing in GPU-based ANNS, but it cannot provide high recall.
To remedy these limitations, we present a novel GPU-based solution wherein we conduct ANNS that does not require the large graph to be present on GPU memory, instead the CPU retrieves the required neighbourhood information from the graph index kept on the host RAM and transfers it to GPU efficiently. Likewise, since the GPU cannot accommodate the entire base dataset, the search process on the GPU employs compressed vectors (instead of the base dataset vectors) for distance calculations. The compressed vectors are generated using the widely utilized Product Quantization [28] technique, employed in numerous previous works, including FAISS [27] and DiskANN [25].
Thus BANG mitigates the CPU-GPU data transfer bottleneck by transferring only compressed data to the GPU and by eliminating the need for transferring the entire graph to the GPU by fetching the neighbour information for a node from CPU to GPU on demand.
@@ III-B Optimal Hardware Usage
BANG’s approach of obtaining neighbours from the CPU and calculating distances from them to the query on the GPU using compressed vectors presents two challenges.
The primary challenge is to prevent CPU and GPU idleness by ensuring their concurrent and continuous utilization despite the dependence on neighbour information between these devices. Furthermore, it is important to consider the volume of data that needs to be transferred between the CPU and GPU when balancing the work distribution between them to ensure that performance is not constrained by bandwidth limitations caused by prolonged data transfers, as in the GGNN [21] framework. To address these issues, we propose a parallel ANNS implementation BANG that extracts maximum parallelism by efficiently utilizing the hardware (CPU, GPU and PCIe bus) with a phased-execution approach to efficiently load balance the ANNS work across CPU and GPU and to minimize CPU-GPU data transfer bottlenecks. The phased-execution strategy facilitates the execution of different steps of ANNS in a pipeline on both CPU and GPU, which enables: (1) CPU-GPU load balancing to minimize CPU and GPU idle time, (2) using prefetching techniques to transfer data from CPU to GPU asynchronously for certain phases, and (3) optimizing the span characteristics of each GPU kernel (corresponding to operations such as distance computation, sorting and updating worklist) separately to maximize parallelism.
The second challenge is to optimize data structure placement according to computational strengths and memory access capabilities of the CPU and GPU, owing to the higher memory footprint caused by thousands of queries running concurrently on the data in the GPU. For instance, when handling parallel queries (Q) on the GPU and managing data structures for tracking visited nodes (V) and their processing, in order to determine whether structures should be placed on the CPU or GPU, several key factors need to be considered, such as the computational complexity associated with the memory size of $Q\times V$, the patterns of memory accesses within V (access times vary with CPU and GPU memory bandwidths), and the time it takes to transfer data $Q\times V$ given the limited bandwidth of the PCIe interconnect between the CPU and GPU. BANG handles this by using an efficient implementation of bloom filter[7] on the GPU. We discuss this in detail in Section IV-C.

### v4

@@ Abstract
Approximate Nearest Neighbour Search (ANNS) is a subroutine in algorithms routinely employed in information retrieval, pattern recognition, data mining, image processing, and beyond. Recent works have established that graph-based ANNS algorithms are practically more efficient than the other methods proposed in the literature. The growing volume and dimensionality of data necessitates designing scalable techniques for ANNS. To this end, the prior art has explored parallelising graph-based ANNS on GPU, leveraging its massive parallelism. The current state-of-the-art GPU-based ANNS algorithms either (i) require both the dataset and the generated graph index to reside entirely in the GPU memory, or (ii) they partition the dataset into small independent shards, each of which can fit in GPU memory, and perform the search on these shards on the GPU. While the first approach fails to handle large datasets due to the limited memory available on the GPU, the latter delivers poor performance on large datasets due to high data traffic over the low-bandwidth PCIe interconnect.
We introduce BANG, a first-of-its-kind technique for graph-based ANNS on GPU for billion-scale datasets, that cannot entirely fit in the GPU memory. BANG stands out by harnessing a compressed form of the dataset on a single GPU to perform distance computations while efficiently accessing the graph index kept on the host memory, enabling efficient ANNS on large graphs within the limited GPU memory. BANG incorporates highly optimised GPU kernels and proceeds in phases that run concurrently on the GPU and CPU, taking advantage of their architectural specificities. Furthermore, it enables overlapping communication with computation that results in efficient data transfer between the CPU and GPU. Using a single NVIDIA Ampere A100 GPU, BANG achieves throughputs 50$\times$-400$\times$ higher than competing methods for a recall of 0.9 on three popular billion-scale datasets.
@@
approximate nearest neighbour search, graph and tree search, information retrieval, approximate search, vector similarity search, GPU, big data.
@@ I Introduction
The $k$-Nearest-Neighbour-Search ($k$NN) problem is to find the $k$ nearest data points to a given query point in a multidimensional dataset. As the dimensionality increases, exact search methods become increasingly inefficient. In order to evaluate the exact $k$-nearest-neighbours of a query point in a $d$-dimensional dataset having $n$ points, as a consequence of the curse of dimensionality, all $n$ points must be examined; this takes $O(nd)$ time [28]. Therefore, it has become commonplace to use Approximate Nearest Neighbour Search (ANNS) for finding the nearest neighbours of a query, to mitigate the curse of dimensionality by sacrificing a small accuracy for speed [68]. ANNS accuracy is typically measured as recall, which quantifies the overlap between the retrieved nearest neighbours and the ground truth for a query. ANNS is widely applied in information retrieval [10, 7, 39], recommendation systems [49, 57], search engines [65], and computer vision [70] to search through large datasets of words, documents, images, and multimedia. Advances in deep learning [6] have made ANNS essential for similarity search [68]. ANNS is a key subcomponent of vector databases [47] and is transforming the way vector embeddings are efficiently indexed and queried [27]. Search queries on such massive multidimensional datasets are often processed in batches to meet the high throughput demands, as evidenced by recent efforts [31, 24, 63, 67] leveraging the massive parallelism of GPUs for ANNS. One of the earliest use cases of GPUs in this context was accelerating $k$NN queries on road networks, such as locating the $k$ nearest cars for ride-sharing [34]. Nowadays, GPUs are increasingly powering vector databases for Retrieval-Augmented Generation (RAG) [45], boosting semantic accuracy and freshness of results in static Large Language Models (LLMs) and supporting high-throughput use cases, such as cloud-based chatbots [5] and unstructured data applications like e-commerce [69].
Billion-scale ANNS is computationally intensive, with pairwise distance calculations well-suited for GPU architectures. Zhang et al. [66] compare CPU- and GPU-based ANNS techniques, and conclude that GPUs offer higher performance and cost-effectiveness for accelerating vector query processing. The Billion-Scale Approximate Nearest Neighbor Search Challenge [53] evaluates ANNS on various hardware platforms (e.g. CPU, GPU, custom hardware) and establishes that GPU-based implementations significantly outperform. Billion-scale datasets exceed single GPU memory, and hence ANNS on these datasets requires multiple GPUs [24]; this introduces two challenges: (a) a significant escalation in implementation cost (as seen in the Track3 Cost/Power Leaderboards [53]), (b) a mismatch between GPU compute performance and memory resources.
@@
Graph-based ANNS algorithms [36, 21, 29] have been shown to be generally more efficient, in practice, at handling large datasets. However, GPU implementations of these algorithms [46, 67, 24] require storing graph data structures in GPU memory, which limits their ability to handle large datasets. Even recent GPUs like the NVIDIA Ampere A100, having 80GB device memory, cannot accommodate the entire input data (graph index and data points). Prior solutions, such as sharding, effectively implemented by GGNN [24], incur high memory transfer cost. For example, with the PCIe 4.0 interconnect operating at its peak theoretical transfer rate of 32 GB/s, transferring the DEEP1B dataset (having a data size of 384GB and a graph index size of 260GB) from CPU to GPU would take 20 seconds. This would result in a low throughput of less than 500 QPS for 10,000 concurrent queries on the GPU. Alternatively, hashing and compression techniques, as showcased by SONG [67] and FAISS [31], can handle large volumes of data using a single GPU by reducing the data dimensionality or by compressing the vectors. However, these approaches face limitations in achieving high recall on massive datasets. It has been shown that it is feasible to attain high throughputs at high recalls on large datasets using multiple GPUs [24]. However, such approaches have a very high hardware cost which makes them undesirable. Thus, this paper explores an important question: Can we increase the throughput of ANNS queries without compromising their recall by using a single GPU?
@@
In this paper, we introduce BANG, a novel GPU-based ANNS method, which is designed to efficiently handle large datasets, with billions of points, that cannot entirely fit in GPU memory. BANG stands out by employing compressed data for accelerated distance computations on the GPU while keeping the graph index and the dataset on the CPU, thus enabling efficient ANNS on large datasets. CPU transfers neighbour information to the GPU on demand in every iteration of the search. Furthermore, in contrast to the previous approaches that treat search activity as one monolithic block, BANG divides ANNS activity into distinct phases, in order to maximise resource utilisation and mitigate CPU-GPU data transfer bottlenecks. Specifically, the division into phases leads to three advantages: (1) executing different phases on CPU and GPU in a pipeline (2) optimising each phase’s (GPU kernel’s) span characteristics separately to maximise parallelism on GPU (3) asynchronously prefetching required data from CPU to GPU for specific phases. Thus, the overall performance of BANG stems from several factors, such as CPU-GPU load balancing, highly optimised GPU kernels (for tasks such as distance calculations, sorting, and updating worklists), prefetching and pipelining.
@@
ANNS methods typically follow the index and search paradigm whereby data points are first processed to construct an efficient index, and search queries are then processed using this index. In this work, our focus is efficient billion-scale ANNS on GPU utilising an underlying graph index, and therefore we do not construct a graph index but instead utilise an existing one from prior work [29].
Our evaluations show that BANG significantly outperforms the state-of-the-art on four popular real-world billion-size ANNS datasets with varying characteristics across all recall values on a single NVIDIA A100 GPU. Notably, on the SIFT-1B and Deep-1B datasets, we achieve 50.5×50.5$\times$50.5 × higher average throughput over competing methods at a high recall of 0.95.
@@
  • We propose a phased-execution strategy that enables pipelined execution of ANNS steps on both CPU and GPU, optimising CPU-GPU load balancing, asynchronous data transfer, and GPU kernel performance for maximising parallelism.
  • We present efficient GPU kernels for distance calculations and worklist updates in ANNS and implement prefetching and pipelining to effectively utilise CPU and GPU, reducing traffic over PCIe interconnect and enhancing throughput while retaining recall.
@@ II-B Graph Index
Several of the best-performing ANNS algorithms are graph-based, achieving both high recall and high throughput, measured as queries per second (QPS). A graph-based ANNS algorithm runs on a proximity graph, which is pre-constructed over the dataset points by connecting each point to its nearby points. Our BANG algorithm implements an efficient ANNS method that can run on various proximity graphs, such as kNN [24], HNSW [36], Vamana [29], and others. Manohar et al. [38] conduct a study on various CPU-based ANNS algorithms and show that the DiskANN [29] approach, which runs on the Vamana graph index, achieves superior throughput and recall on large datasets. Furthermore, the recent work by Wang et al. [59], which proposes enhancements to the Vamana graph index for disk-resident indexes, also highlights the significance of the Vamana graph. Hence, we employ the Vamana graph as the underlying graph index in the implementation of BANG. Vamana graph index is constructed by sharding the large dataset into smaller overlapping clusters. The technique iteratively processes individual clusters and uses RobustPrune and GreedySearch routines to construct directed subgraphs on the host RAM, which are finally written to an SSD. The final graph is formed by merging the edges in the subgraphs. While prior techniques that construct the graph index using the SNG property [4] aim to reduce distances to the query point along the search path, Vamana further reduces the number of hops to the query point with the RobustPrune property, resulting in a faster search.
@@ III ANN Search on GPU: Challenges in Handling Billion-scale Data
We propose BANG, a method for parallel ANNS on a single GPU. There are two main challenges in parallelising ANNS on a single GPU: (1) managing large graphs within the constraints of limited GPU memory, (2) extracting sufficient parallelism for ANNS to fully utilise the hardware resources. In the following subsections, we discuss these challenges in detail.
@@ III-A Limited GPU Memory
A primary challenge when dealing with large datasets on GPUs is their massive memory footprint. Table I shows the sizes of various billion-scale ANNS datasets and their respective graph sizes. Notably, even with the recent A100 GPU’s maximum global memory capacity of 80GB, it is evident that the entire input data (base, graph, and query) cannot be accommodated within the GPU memory. To address this challenge, a recent work GGNN [24] effectively implemented sharding as a solution. However, this approach requires frequent swaps of both processed and unprocessed shards between GPU and host RAM, resulting in high memory transfer costs over the PCI interconnect. Further, these swaps may result in an idle time of CPU/GPU, and optimal GPU-CPU coordination is essential during iterative graph traversals of ANNS. Furthermore, sharding the entire data is not viable because even with the PCIe 4.0 interconnect operating at its peak theoretical transfer bandwidth of 32 GB/s, it will take an estimated 20 seconds to transfer the largest dataset from CPU to GPU and will result in a markedly low throughput of less than 500 QPS as explained in Section I.
Although employing a multi-GPU setup can facilitate the effective distribution of shards across all GPUs to store the entire input data (as demonstrated with eight GPUs in GGNN [24]), this approach incurs significant hardware cost. As a result, our focus is on developing an efficient and cost-effective parallel implementation of ANNS using a single GPU.
@@
SIFT-1B 128 GB 260 GB 80 GB
Deep-1B 384 GB 260 GB 96 GB
MSSPACEV-1B 100 GB 260 GB 80 GB
Text-to-Image-1B 800 GB 260 GB 80 GB
Alternatively, hashing and quantization techniques can handle large data on a single GPU. Hashing effectively reduces data dimensionality to make it fit within GPU memory, as demonstrated by SONG [67], which compresses a 784-byte vector to 64 bytes in the MNIST8M dataset. However, hashing is better suited for smaller datasets and may not achieve high recall on datasets with billions of points. Data compression, as demonstrated by FAISS [31], can accelerate query processing in GPU-based ANNS, but it cannot provide high recall.
To mitigate these limitations, we present a novel GPU-based solution wherein we conduct ANNS that does not require the large graph to be present in GPU memory. Instead, the CPU retrieves the required neighbourhood information from the graph index kept on the host RAM and transfers it to the GPU efficiently. Likewise, since the GPU cannot accommodate the entire base dataset, the search process on the GPU employs compressed vectors (instead of the base dataset vectors) for distance calculations. The compressed vectors are generated using the widely utilised Product Quantization [32] technique, employed in several previous works, including FAISS [31] and DiskANN [29]. Thus, BANG alleviates the CPU-GPU data transfer bottleneck by transferring only compressed data to the GPU and eliminating the need to transfer the entire graph. Instead, it fetches neighbour information for a node from the CPU to the GPU on demand.
@@ III-B Optimal Hardware Usage
BANG’s approach of fetching neighbours from the CPU and calculating distances from them to the query on the GPU using compressed vectors presents two challenges.
The primary challenge is to prevent CPU and GPU idleness by ensuring their concurrent and continuous utilisation despite the dependence on neighbour information between these devices. Furthermore, it is important to consider the volume of data that needs to be transferred between the CPU and the GPU when balancing the work distribution between them to ensure that performance is not constrained by bandwidth limitations caused by prolonged data transfers, as in the GGNN [24] framework. To address these issues, we propose a parallel ANNS implementation BANG that maximises parallelism by efficiently utilising the hardware (CPU, GPU and PCIe bus) with a phased-execution approach to efficiently load balance the ANNS work across CPU and GPU, and to minimise CPU-GPU data transfer bottlenecks. The phased-execution strategy facilitates the execution of different steps of ANNS in a pipeline on both CPU and GPU, which enables: (1) CPU-GPU load balancing to minimise CPU and GPU idle time, (2) using prefetching techniques to transfer data from CPU to GPU asynchronously for certain phases, and (3) optimising the span characteristics of each GPU kernel (corresponding to operations such as distance computation, sorting and updating worklist) separately to maximise parallelism.
The second challenge is to optimise the placement of the data structure according to the computational strengths and memory access capabilities of the CPU and GPU, due to the higher memory footprint caused by thousands of queries running concurrently on the data in the GPU. For example, when handling parallel queries (Q) on the GPU and managing data structures for tracking visited nodes (V) and their processing, in order to determine whether structures should be placed on the CPU or GPU, several key factors need to be considered, such as the computational complexity associated with the memory size of $Q\times V$, the patterns of memory accesses within V (access times vary with CPU and GPU memory bandwidths), and the time it takes to transfer data $Q\times V$ given the limited bandwidth of the PCIe interconnect between the CPU and GPU. BANG handles this by using an efficient implementation of bloom filter [8] on the GPU. We discuss this in detail in Section IV-C.